# 3DCoMPaT++: An improved Large-scale 3D Vision Dataset for Compositional Recognition

### v2

@@ 4.2 Grounded Compositional Recognition (GCR)
Fine-grained PointNet+SegFormer [28, 29] 84.18 42.57 9.05 26.68 3.84
@@
Coarse-grained PointNet+SegFormer [28, 29] 84.27 65.61 44.82 52.82 29.74
@@
TABLE VI: Grounded Compositional Recognition (GCR). We evaluate the performance of various baselines on the GCR task. While modality fusion-based methods like BPNet [34] and PointNet+SegFormer [28, 29] perform well on the shape classification task they still underperformed compared to the RGB pointcloud-based baseline. We report the GCR metrics under both fine-grained and coarse-grained settings, using 10 compositions per shape. Overall, recognizing and grouding all part-material pairs of a shape is particularly challenging, especially in the fine-grained setting.
  • Value. Proportion of correctly predicted part-material pairs.
We consider a part to be correctly segmented if the predicted part segmentation mask has an intersection over union (IoU) of at least 0.5 with the ground-truth part segmentation mask. In 2D, we use the pixel-wise definition of IoU. For the 3D modality, we use the point-wise definition.
Note that Value and Grounded-value are both evaluated at the shape level: we divide the number of correctly identified (resp. grounded) part-material pairs by the total number of parts appearing in each shape, and then average across all samples. Value is thus upper bounded by Value-all, and Grounded-value by Grounded-value-all.
@@
  • “PointNet+SegFormer”. This baseline employs separate 2D/3D models and fuses predictions at evaluation time. We use PointNeXT [30] for 3D shape classification and SegFormer [29] for 2D material segmentation and 2D part segmentation. 2D dense predictions are projected to the 3D space using the depth maps and camera parameters. We use this baseline to assess the feasibility of the GCR task on 3DCoMPaT++ when all part-pair predictions are performed on the 2D space.
@@
Results. Table VI summarizes GCR results of baseline methods and challenge winners. The PointNet+SegFormer 2D-based baseline is markedly outperformed by the BPNet multimodal baseline, which can be imputed to the absence of explicit 3D-aware modality fusion during training. The winning method PointNet++RGB, which takes only 3D point clouds as inputs and leverages a powerful point grouping module, performs best in the coarse-grained setting, reaching 58.99% on the Grounded-value-all metric. BPNet on the other hand, performs best in the fine-grained setting, and outperforms PointNet++RGB on the classification metrics (Value and Value-all), but is overtaken by PointNet++RGB on the grounded metrics. More importantly, we notice that even our best performing models struggle to perform on the fine-grained GCR task, where we reach most 17.55% on the Grounded-value-all metric. This suggests that creating a single model able to achieve strong performance across GCR metrics poses great challenges, especially in the fine-grained setting.
@@
5 Conclusion
We introduce 3DCoMPaT++, a large-scale dataset of Compositions of Materials on Parts of 3D Things, which contains 10M styled models stemming from 10000 3D shapes from 41 object categories. 3DCoMPaT++ contains 3D shapes, part segmentation information in fine-grained and coarse-grained semantic levels and material compatibility information, so that multiple high-quality PBR materials can be assigned to the same shape part. We also propose a new task, dubbed as Grounded CoMPaT Recognition (GCR), that our dataset enables and introduces baseline methods to solve them. Future directions may include additional tasks such as 3D part-aware shape synthesis, 3D part-aware reconstruction from 2D views, and 3D part-style transfer, which all can be enabled by the rich data provided in 3DCoMPaT++.
@@ Acknowledgements
Figure 16: Randomly sampled styles for a set of six geometries from the 3DCoMPaT++ dataset. Stylized shapes are sampled with a random style parameterized by a set of compatible materials.
Figure 17: Canonical (first four columns) and random views (last four columns) rendered for various stylized shapes and shape categories. Random viewpoints are sampled, while canonical viewpoints are equally spaced around the shape.
Figure 18: Additional examples for data provided for six diverse shapes, following the first canonical viewpoint. Each 2D image is complemented by corresponding part segmentation masks, material masks, and absolute depth maps. Camera parameters are also provided for each render. For each stylized shape, we also provide RGB pointclouds labeled at the part and material levels. Part segmentation data is available in both coarse-grained and fine-grained resolutions.
Appendix D
@@
Figure 19: Going from fine-grained to coarse-grained segmentation. We consider shapes from the ”cabinet”, ”bag” and ”faucet” categories. Fine-grained classes are merged following a shape category-specific nomenclature to create coarse-grained classes. Resulting shapes are simplified and contain fewer parts.
Appendix E
@@
Figure 20: Excerpts from the 3DCoMPaT++ guidelines. We show excerpts from the 3DCoMPaT++ guidelines for the faucet, shower, sink shape categories. Related shape categories with a large number of shared parts are defined jointly in the guidelines. In total, 33 annotation guidelines were created and refined for the 41 shape categories of 3DCoMPaT++, for a total of 606 pages of documentation.

### v3

@@ IV-B Grounded Compositional Recognition (GCR)
TABLE VI: Grounded Compositional Recognition (GCR). We evaluate the performance of various baselines on the GCR task. While modality fusion-based methods like BPNet [40] and PointNeXt+SegFormer [36, 30] perform well on the shape classification task they still underperformed compared to the RGB pointcloud-based baseline. We report the GCR metrics under both fine-grained and coarse-grained settings, using 10 compositions per shape.
@@
Fine-grained PointNeXt+SegFormer [36, 30] 84.18 42.57 9.05 26.68 3.84
@@
Coarse-grained PointNeXt+SegFormer [36, 30] 84.27 65.61 44.82 52.82 29.74
@@
  • “PointNeXt+SegFormer”. This baseline employs separate 2D/3D models and fuses predictions at evaluation time. We use PointNeXt [36] for 3D shape classification and SegFormer [30] for 2D material segmentation and 2D part segmentation. 2D dense predictions are projected to the 3D space using the depth maps and camera parameters. We use this baseline to assess the feasibility of the GCR task on 3DCoMPaT++ when all part-pair predictions are performed on the 2D space.
@@
Results. Table VI summarizes GCR results of baseline methods and challenge winners. The PointNeXt+SegFormer 2D-based baseline is markedly outperformed by the BPNet multimodal baseline, which can be imputed to the absence of explicit 3D-aware modality fusion during training. The winning method PointNet++RGB, which takes only 3D point clouds as inputs and leverages a powerful point grouping module, performs best in the coarse-grained setting, reaching 58.99% on the Grounded-value-all metric. BPNet on the other hand, performs best in the fine-grained setting, and outperforms PointNet++RGB on the classification metrics (Value and Value-all), but is overtaken by PointNet++RGB on the grounded metrics. More importantly, we notice that even our best performing models struggle to perform on the fine-grained GCR task, where we reach at most 17.55% accuracy on the Grounded-value-all metric. Overall, recognizing and grouding all part-material pairs of a shape is particularly challenging, especially in the fine-grained setting.
@@
V Conclusion
We introduce 3DCoMPaT++, a large-scale dataset of Compositions of Materials on Parts of 3D Things, which contains 10M styled models stemming from 10000 3D shapes from 42 object categories. 3DCoMPaT++ contains 3D shapes, part segmentation information in fine-grained and coarse-grained semantic levels and material compatibility information, so that multiple high-quality PBR materials can be assigned to the same shape part. We also propose a new task, dubbed as Grounded CoMPaT Recognition (GCR), that our dataset enables and introduces baseline methods to solve them. Future directions may include additional tasks such as 3D part-aware shape synthesis, 3D part-aware reconstruction from 2D views, and 3D part-style transfer, which all can be enabled by the rich data provided in 3DCoMPaT++.
@@ Acknowledgements
Additional mesh quality statistics compared to the original ShapeNet [1] dataset and the ABO [5] dataset are provided in Figure 16, including the number of vertices, faces, and edges per shape, percentage of watertight meshes and percentage of meshes with degenerate faces.
Appendix B
Rendered representative shapes from every 42 shape category in 3DCoMPaT++ in Figure 17.
Appendix C
Rendered materials from every 293 material category in 3DCoMPaT++ in Figure 18.
Appendix D
@@
Figure 16: Mesh quality statistics. Statistics on the mesh quality of 3DCoMPaT++ compared to the original ShapeNet [1] dataset, and the ABO [5] dataset.
Top: Number of vertices, faces and edges per 3D shape. Bottom: Percentage of meshes with non-manifold geometries and percentage of meshes with degenerate meshes. Overall, we observe that 3DCoMPaT++ has a higher percentage of watertight meshes and lower percentage of degenerate meshes compared to the original ShapeNet dataset. Our dataset also compares favorably to the ABO dataset, which contains almost only meshes with non-manifold geometries.
Figure 17: Rendered shapes from 3DCoMPaT++. We render representative shapes from every 42 shape category in 3DCoMPaT++, with the shape category indexes and names.
Figure 18: Rendered materials. We render the material collection we use from every part-material assignment in our dataset. We display 293 fine-grained material categories, grouped by coarse material category (continued in the next page).
Figure 19: Generated styles. Randomly sampled styles for a set of six geometries from the 3DCoMPaT++ dataset. Styles are generated by sampling a set of compatible materials from the material collection, using the style generation algorithm described in Figure 13.
Figure 20: Canonical and random views. Canonical (first four columns) and random views (last four columns) rendered for various stylized shapes and shape categories. Random viewpoints are sampled, while canonical viewpoints are equally spaced around the shape.
Figure 21: Complete data samples. Additional examples for data provided for six diverse shapes, following the first canonical viewpoint. Each 2D image is complemented by corresponding part segmentation masks, material masks, and absolute depth maps. Camera parameters are also provided for each render. For each stylized shape, we also provide RGB pointclouds labeled at the part and material levels. Part segmentation data is available in both coarse-grained and fine-grained resolutions.
Appendix G
@@
Figure 22: From fine-grained to coarse-grained segmentation. We consider shapes from the ”cabinet”, ”bag” and ”faucet” categories. Fine-grained classes are merged following a shape category-specific nomenclature to create coarse-grained classes. Resulting shapes are simplified and contain fewer parts.
Appendix H
@@
Figure 23: Excerpts from the 3DCoMPaT++ guidelines. We show excerpts from the 3DCoMPaT++ guidelines for the faucet, shower, sink shape categories. Related shape categories with a large number of shared parts are defined jointly in the guidelines. In total, 33 annotation guidelines were created and refined for the 42 shape categories of 3DCoMPaT++, for a total of 606 pages of documentation.